OpenStudy (anonymous):

I need some help with rational inequalites it's kicking mu butt. 4x+7/-3 <= 2x+5

OpenStudy (anonymous):

assuming:(4x+7)/-3 <= 2x+5 You get x >= -2.2 Kee in mind the inequality flips when you multiply both sides by a negative no. Hope I was of some help

OpenStudy (anonymous):

I need to complete this using algebra, I think I'm completely messing up the +.- signs

OpenStudy (anonymous):

\[(4x + 7)/ (-3) \le 2x + 5 \] => \[4x + 7 \ge -3 ( 2x + 5)\] =>\[4x + 7 \ge -6x -15\] =>\[10x \ge -22\] => \[x \ge -2.2\]